# Reliable Wireless Indoor Localization via Cross-Validated Prediction-Powered Calibration

### v2

@@ I Introduction
When unlabeled data are available, one can construct a synthetic dataset with pseudo-labels generated by an existing predictive model. Prediction-powered inference (PPI) [10] is a recently proposed framework for incorporating model-generated pseudo-labels from an auxiliary predictor, while preserving statistical validity. While PPI applies to parameter estimation, the work [11] adapted PPI as a mechanism to construct prediction sets via RCPS using both real and synthetic labels—an approach referred to as henceforth as RCPS-PPI.
Figure 1: Comparison of risk-controlling prediction sets for the task of wireless indoor localization of mobile devices under three calibration strategies: RCPS [9], which uses only real-world labeled data; RCPS-PPI [11], which splits the real data into two subsets—one for fine-tuning the label-generating predictive model and the other for estimating the model-induced bias on synthetic labels; and the proposed RCPS-CPPI, which uses the entire labeled dataset for both predictor fine-tuning and bias estimation via cross-validation, yielding more efficient prediction sets without compromising coverage.
The key challenge in PPI—and in the PPI-based RCPS approach in [11]—is that the pseudo-labels are generally biased estimates of the true labels. This may violate statistical validity if used naively. PPI addresses this problem by applying a bias correction using a small labeled dataset. However, a limitation of PPI is that it requires splitting the labeled data. In fact, a portion of the dataset must be set aside to train or fine-tune the auxiliary label-generating predictor on the given inference task, while the remaining portion of the labeled dataset is used for bias correction. Given limited labeled data, such splitting is inefficient and can degrade performance.
Cross-PPI (CPPI) [12] was recently proposed to overcome this issue. CPPI uses $K$-fold cross-validation to utilize all labeled samples for both predictor training and calibration [13]. However, the use of CPPI is currently limited to parameter estimation.
In this letter, we develop a calibration scheme that integrates CPPI with the RCPS framework to enhance the efficiency of prediction sets by leveraging synthetic pseudo-labels. The proposed method, termed RCPS-CPPI, uses the available labeled data to simultaneously fine-tune a predictor and estimate the bias of the pseudo-labels, yielding prediction sets that satisfy a target risk level with a user-defined confidence probability.
As illustrated conceptually in Fig. 1, RCPS-CPPI produces tighter prediction sets than existing ones that do not leverage synthetic labels or apply PPI as in [11]. While RCPS can be applied to any inference task, we validate it on an indoor localization task [14], demonstrating valid coverage with significantly reduced prediction set size compared to baseline approaches.
@@ II-B Risk-Controlling Prediction Sets
Specifically, we aim at augmenting a decision $f(X)$ for any input $X$ with a prediction set $\Gamma_{\lambda}(X)\subseteq\mathcal{Y}$, depending on a threshold parameter $\lambda$, that satisfies given statistical guarantees. The statistical performance of the set $\Gamma_{\lambda}(X)\subseteq\mathcal{Y}$ is measured by a loss function $\ell(Y,\Gamma_{\lambda}(X))$, such as the miscoverage loss
@@
where $\mathbbm{1}\{\cdot\}$ is the indicator function. In the example of Fig. 1, this loss measures whether the prediction set $\Gamma_{\lambda}(X)$ includes the true UE location $Y$, yielding $\ell(Y,\Gamma_{\lambda}(X))=0$, or not, producing $\ell(Y,\Gamma_{\lambda}(X))=1$.
Formally, the target statistical requirement is that the expected risk
@@
where the expectation is over the distribution $P_{XY}$ of the test data $(X,Y)$, is no larger than a threshold $\alpha$ with probability at least $1-\delta$, i.e.,
@@
Following prior art [9, 11], we make the following technical assumptions. The loss function $\ell(Y,\Gamma_{\lambda}(X))$ is bounded between 0 and 1, and is non-increasing as the prediction set grows—i.e., for $\lambda^{\prime}\leq\lambda$, and hence for $\Gamma_{\lambda}(X)\supseteq\Gamma_{\lambda^{\prime}}(X)$, we have the inequality $\ell(Y,\Gamma_{\lambda}(X))\leq\ell(Y,\Gamma_{\lambda^{\prime}}(X))\,$. These conditions are satisfied by the miscoverage loss (1).
@@ III-A Risk-Controlling Prediction Sets based on Real Data
where the probability is over the calibration dataset $\mathcal{D}^{\text{L}}$. The UCB can be constructed by leveraging the boundedness of the loss via methods such as the Waudby-Smith-Ramdas (WSR) estimator [17]. Finally, RCPS choose the threshold as
@@ IV-B RCPS-CPPI
and applies selection rule (8) using the estimate $\hat{R}^{+}_{\text{CPPI}}(\lambda)$ instead of $\hat{R}^{+}(\lambda)$.
@@ V-A Setup
We evaluate the proposed approach on a wireless indoor localization task using a WiFi fingerprinting dataset [14, 18]. We randomly sample a subset of 100 labeled examples to train the base model $f$ and simulate scenarios with limited calibration datasets $\mathcal{D}^{\text{L}}$, with $n$ varying from 50 up to 500 labeled calibration points. The remaining data are used as an unlabeled calibration dataset $\mathcal{D}^{\text{U}}$. We reserve 30 labeled samples for a test set to evaluate coverage and set size.
@@ V-B Results
We report the empirical coverage, i.e., the fraction of test points whose true location lies inside the prediction set, and the inefficiency, defined as the average radius of the prediction sets $\Gamma_{\hat{\lambda}}(X)$, on the test samples. Apart from RCPS (Section III-A), and RCPS-PPI (Section III-B), we also consider a baseline semi-supervised (SS) scheme that uses both labeled and unlabeled data without any bias correction and directly applies RCPS on the combined data.

### v3

@@ I Introduction
When unlabeled data are available, one can construct a synthetic dataset with pseudo-labels generated by an existing predictive model. Prediction-powered inference (PPI) [2] is a recently proposed framework for incorporating model-generated pseudo-labels from an auxiliary predictor, while preserving statistical validity. While PPI applies to parameter estimation or learning, the work [5] adapted PPI as a mechanism to construct prediction sets via RCPS using both real and synthetic labels—an approach referred to henceforth as RCPS-PPI.
Figure 1: Comparison of risk-controlling prediction sets (concentric circles) for the task of wireless indoor localization of mobile devices under three calibration strategies: RCPS [3], which uses only real-world labeled data; RCPS-PPI [5], which splits the real data into two subsets for fine-tuning the label-generating predictive model and for estimating the bias caused by synthetic labels; and the proposed RCPS-CPPI, which uses the entire labeled dataset for both predictor fine-tuning and bias estimation via cross-validation (see Table I).
TABLE I: Comparison of state-of-the-art schemes.
Scheme Prediction / Learning Synthetic data Cross-validation
RCPS [3] Prediction
PPI [2] Learning
RCPS-PPI [5] Prediction
CPPI [19], [16] Learning
XB [8] Prediction
RCPS-CPPI Prediction
In RCPS-PPI, a portion of the dataset must be set aside to train or fine-tune the auxiliary label-generating predictor on the given inference task, while the remaining portion of the labeled dataset is used for bias correction. Given limited labeled data, such splitting is inefficient and can degrade performance.
Cross-PPI (CPPI) [19] was recently proposed to overcome this issue. CPPI uses $K$-fold cross-validation to utilize all labeled samples for both predictor training and calibration. However, while the use of CPPI is currently limited to learning, cross-validation was applied to set prediction in [8]. A summary of the state of the art can be found in Table I.
In this letter, we develop a calibration scheme that enhances the efficiency of prediction sets by leveraging synthetic pseudo-labels. The proposed method, termed RCPS-CPPI, uses the available labeled data to simultaneously fine-tune a predictor and estimate the bias of the pseudo-labels, yielding prediction sets that satisfy a target risk level with a user-defined confidence probability. While RCPS can be applied to any inference task, we validate it on an indoor localization task [10], demonstrating valid coverage with significantly reduced prediction set size compared to baseline approaches.
@@ II-B Risk-Controlling Prediction Sets
Specifically, we aim at augmenting a decision $f(X)$ for any input $X$ with a prediction set $\Gamma_{\lambda}(X)\subseteq\mathcal{Y}$, depending on a threshold parameter $\lambda\in[0,\lambda_{\max}]$, that satisfies given statistical guarantees. The statistical performance of the set $\Gamma_{\lambda}(X)\subseteq\mathcal{Y}$ is measured by a loss function $\ell(Y,\Gamma_{\lambda}(X))$, such as the miscoverage loss
@@
where $\mathbbm{1}\{\cdot\}$ is the indicator function. In the example of Fig. 1, this loss provides a direct measure of the performance of the prediction set for localization by returning $\ell(Y,\Gamma_{\lambda}(X))=0$ if the prediction set $\Gamma_{\lambda}(X)$ includes the true UE location $Y$, while yielding $\ell(Y,\Gamma_{\lambda}(X))=1$ otherwise.
This way the expected risk
@@
where the expectation is over the distribution $P_{XY}$ of the test data $(X,Y)$, measures the probability that the prediction set fails to include the true UE location. The target statistical requirement is that this expected risk is no larger than a threshold $\alpha$ with probability at least $1-\delta$, i.e.,
@@
Following prior art [3, 5], we make the following technical assumptions. The loss function $\ell(Y,\Gamma_{\lambda}(X))$ is bounded between 0 and 1; is non-increasing as the prediction set grows—i.e., for $\lambda^{\prime}\leq\lambda$, and hence for $\Gamma_{\lambda}(X)\supseteq\Gamma_{\lambda^{\prime}}(X)$, we have the inequality $\ell(Y,\Gamma_{\lambda}(X))\leq\ell(Y,\Gamma_{\lambda^{\prime}}(X))\,$; and we have $\ell\big(Y,\Gamma_{\lambda_{\max}}(X)\big)=0$. These assumptions are satisfied by the miscoverage loss (1), as well as other standard losses, including the false negative rate in multi-label classification [3].
@@ III-A Risk-Controlling Prediction Sets based on Real Data
where the probability is over the calibration dataset $\mathcal{D}^{\text{L}}$. The UCB can be constructed by leveraging the boundedness of the loss via methods such as the Waudby-Smith-Ramdas (WSR) estimator [12]. Finally, RCPS chooses the threshold as
@@ IV-B RCPS-CPPI
and applies selection rule (8) using the estimate $\hat{R}^{+}_{\text{CPPI}}(\lambda)$ instead of $\hat{R}^{+}(\lambda)$. Using grid search over parameter $\lambda$ with $m$ points, the complexity of this operation is of order $\mathcal{O}(K(N+n)m)$.
@@ V-A Setup
We evaluate the proposed approach on a wireless indoor localization task using a WiFi fingerprinting dataset [10, 4]. We randomly sample a subset of 100 labeled examples to train the base model $f$ and simulate scenarios with limited calibration datasets $\mathcal{D}^{\text{L}}$, with $n$ varying from 50 up to 500 labeled calibration points. The remaining data are used as an unlabeled calibration dataset $\mathcal{D}^{\text{U}}$. We reserve 30 labeled samples for a test set to evaluate coverage and set size [15].
@@ V-B Results
We report the empirical coverage, i.e., the fraction of test points whose true location lies inside the prediction set, and the inefficiency, defined as the average radius of the prediction sets $\Gamma_{\hat{\lambda}}(X)$, on the test samples. A larger inefficiency implies a more significant uncertainty on the position of the UE, which in turn affects the performance of location-based functionalities such as beamforming and resource allocation [11]. Apart from RCPS (Section III-A), and RCPS-PPI (Section III-B), we also consider a baseline semi-supervised (SS) scheme that uses both labeled and unlabeled data without any bias correction and directly applies RCPS on the combined data.